# Static for Dynamic: Towards a Deeper Understanding of Dynamic Facial Expressions Using Static Expression Data

### v2

@@ IV-D6 Ablation Study on the Position of MoAE
To identify the optimal placement of the MoAE module within the ViT architecture, we conducted ablation studies at different positions. Table VIII shows that placing the MoAE in the last six layers (Later) yields the highest UAR/WAR performance at 43.41%/53.65% on FERV39K and 70.58%/79.37% on DFEW, respectively. In contrast, placing the MoAE in the first six layers (Early) results in a lower UAR/WAR of 43.15%/53.46% on FERV39K and 67.93%/78.42% on DFEW, suggesting that early placement may hinder the model’s ability to learn generic features. Middle and alternate layer placements produce intermediate results. These findings indicate that the MoAE module is most effective in the later layers, where it can enhance task-specific learning by leveraging the learned generic representations.
@@
IV-D7 Ablation Studies on Hyperparameters
Our experiments first focused on the structural parameters of the MoAE module. The results indicate that the model achieves optimal performance in both the FERV39K and DFEW datasets when the MoAE module is configured with eight experts, as illustrated in Fig. 3(a). Any deviation from this optimal number, whether an increase or decrease, results in marginal performance degradation. Similarly, our investigation into the impact of MoAE layer count reveals that incorporating six MoAE layers yields the best performance, as shown in Fig. 3(b). Deviating from this number leads to a noticeable decline in performance. These findings suggest that replacing the final six ViT layers with MoAE layers establishes an optimal balance between task-specific adaptation and preservation of general knowledge.
Beyond structural parameters, we also explored the impact of external training data. As depicted in Fig. 3(c), increasing the proportion of the SFER dataset during the pre-training phase leads to consistent performance improvements on the DFEW dataset (up to 2.87%/3.04% UAR/WAR gain), whereas the impact on FERV39k is less pronounced (approximately 0.49%/1.07% UAR/WAR improvement). These differences may reflect variations in dataset characteristics, such as scale, label noise, or inherent task difficulty. Both datasets achieve peak performance when the entire SFER dataset is employed during pre-training, indicating that pre-training on SFER data is generally advantageous, especially for DFEW. Regarding the SFER dataset utilization during fine-tuning, our experiments reveal that employing 50% of the SFER data per epoch produces optimal results, as depicted in Fig. 3(d). However, increasing the proportion beyond 50% offers no additional benefits and may potentially bias the optimization trajectory of the DFER task.
Overall, these empirical findings highlight the importance of careful hyperparameter tuning in maximizing the effectiveness of the proposed unified learning framework. Specifically, using eight MoAE experts, six MoAE layers, and the entire SFER dataset during pre-training, along with partial SFER data during fine-tuning, establishes a robust foundation for practical deployment while maintaining computational efficiency.
@@ IV-D7 Ablation Studies on Hyperparameters
IV-E Correlation Analysis between SFER and DFER
IV-E1 Evaluation on Cross-Task
Since S4D can handle both SFER and DFER tasks using a shared backbone and respective classifiers, we treat it as a distinct task model during inference. To investigate the correlations and differences between the SFER and DFER tasks, we conducted cross-task evaluation comparisons on the AffectNet-7, FERV39K, and DFEW datasets. As shown in Table IX, the SFER model trained on AffectNet-7 achieves 32.34%/35.10% and 43.36%/47.13% UAR/WAR on the FERV39K and DFEW datasets, respectively. Although this performance is lower than that of DFER models on their respective tasks, it significantly exceeds random guessing (approximately 14.3%), indicating that the SFER model captures relevant spatial features from video data. In contrast, DFER models trained on the FERV39K and DFEW datasets achieve 51.46%/51.38% and 45.20%/45.04% UAR/WAR on AffectNet-7, respectively. These results suggest that DFER models can leverage static expression information from dynamic data, transferring well to the SFER task. We attribute this phenomenon to the inherent presence of static facial expressions within dynamic sequences, which capture peak expressions and serve as snapshots within the temporal progression of the video. However, direct comparisons of UAR/WAR metrics between SFER and DFER models are not ideal due to the inherent differences between the tasks. Therefore, these results should be viewed as reflecting feature transferability rather than a direct performance comparison.
@@ IV-E1 Evaluation on Cross-Task
TABLE IX: Cross-task evaluation comparisons on SFER (AffectNet-7) and DFER (FERV39K, DFEW).
@@
IV-E2 Semantic Relevance between SFER and DFER Tasks
@@
Figure 5: Visualization of the activation distribution of experts in MoAE layers: FERV39K vs. AffectNet-7 and DFEW vs. AffectNet-7 datasets.
@@
IV-E3 Distribution of the Expert Activation in MoAE Layers
Figs. 5(a) and 5(b) show the various activation patterns of experts within the MoAE layers: FERV39K (DFER) vs. AffectNet-7 (SFER) and DFEW (DFER) vs. AffectNet-7 (SFER) datasets. In the FERV39K dataset, expert E3 demonstrates peak activation in layer L4, indicating its specialized capability in processing dynamic facial expressions. Conversely, for AffectNet-7, expert E7 shows pronounced activation in layer L5, suggesting its particular aptitude for SFER. This pattern of task-specific specialization is further validated by the activation patterns observed between DFEW and AffectNet-7 datasets. Specifically, expert E3 exhibits strong activation in layer L4 when processing DFEW data while maintaining relatively lower activation levels for AffectNet-7. Similarly, expert E6 displays higher activation in layer L5 for AffectNet-7, with comparatively reduced activation levels for DFEW. Notably, experts like E1 and E2 display moderate activation across multiple layers in both datasets, suggesting their role in capturing common features shared between tasks. These patterns reveal that while MoAE captures shared representational knowledge between SFER and DFER tasks, it effectively distinguishes task-specific characteristics through expert specialization across layers. Additionally, the similar expert activation patterns observed across different experiments for the AffectNet-7 dataset can be attributed to the same initialization conditions, but also hint at the potential formation of specialized processing pathways for SFER tasks.
@@ IV-F2 Confusion Matrix Analysis
We performed a comprehensive statistical analysis of S4D’s performance on the DFEW [11], FERV39k [9], and MAFW [10] datasets, with the results presented through confusion matrices. As depicted in Fig. 7, major categories such as happy, sad, and neutral expressions demonstrate high accuracy across all datasets. In contrast, minor categories, including disgust, fear, and helplessness exhibit lower accuracy, particularly on the DFEW dataset. This discrepancy likely arises from the limited sample sizes and data imbalance within these categories. Such challenges can be alleviated by implementing imbalanced learning strategies, such as oversampling techniques. Additionally, we observed that some expressions are frequently misclassified as neutral, particularly disgust and surprise, due to the subtle boundaries between neutral and these emotional expressions. For example, in both the DFEW and FERV39K datasets, disgust is frequently misclassified as neutral, likely because of the subtlety of disgust expressions, such as slight nose wrinkling or minor lip movements.

### v3

@@ IV-D6 Ablation Study on the Position of MoAE
To identify the optimal placement of the MoAE module within the ViT architecture, we conducted ablation studies at different positions. Table VIII shows that placing the MoAE in the last six layers (Later) yields the highest UAR/WAR performance at 43.41%/53.65% on FERV39K and 70.58%/79.37% on DFEW, respectively. In contrast, placing the MoAE in the first six layers (Early) results in a lower UAR/WAR of 43.15%/53.46% on FERV39K and 67.93%/78.42% on DFEW, suggesting that early placement may hinder the model’s ability to learn generic features. Middle and alternate layer placements produce intermediate results. These findings indicate that the MoAE module is most effective in the later layers, where it can enhance task-specific learning by leveraging the learned generic representations. Conversely, placing them too early could disrupt the formation of these foundational generic features.
@@
TABLE IX: Ablation Studies on the Masking Ratios. AFC: AffectNet; VC2: VoxCeleb2
Masking Ratio FERV39K DFEW
AFC VC2 UAR WAR UAR WAR
75% 75% 42.03 52.50 65.43 76.97
75% 90% 43.43 52.67 65.95 77.27
75% 95% 42.54 52.47 66.65 77.35
90% 75% 42.28 52.64 68.51 78.25
90% 90% 42.82 52.96 68.10 78.08
90% 95% 43.41 53.65 70.58 79.37
IV-D7 Ablation Study on the Masking Ratios
To assess the impact of masking ratios, we perform an ablation study during pre-training on AffectNet (images) and VoxCeleb2 (videos). As presented in Table IX, the combination of 90% image and 95% video masking ratios consistently yields the best performance, achieving 43.41%/53.65% UAR/WAR on FERV39K and 70.58%/79.37% on DFEW. Notably, this high masking ratio significantly reduces the computational cost during pre-training. Specifically, it reduces FLOPs by approximately 42% for image inputs compared to the MAE [18] with a 75% masking ratio, and by 22% for video inputs compared to VideoMAE [8] with a 90% masking ratio. These findings indicate that an aggressive masking strategy is not only effective in performance but also computationally efficient for fine-grained tasks such like DFER. Based on this analysis, we adopt a 90% masking ratio for images and a 95% masking ratio for videos in our final pre-training models.
@@ IV-D7 Ablation Study on the Masking Ratios
IV-D8 Ablation Studies on Hyperparameters
Our experiments first focused on the structural parameters of the MoAE module. The results indicate that the model achieves optimal performance in both the FERV39K and DFEW datasets when the MoAE module is configured with 8 experts, as illustrated in Fig. 3(a). Any deviation from this optimal number, whether an increase or decrease, results in marginal performance degradation. This suggests that 8 experts strike an effective balance between sufficient model capacity to learn diverse features and mitigating the risks of overfitting or excessive computational overhead that might arise from a larger number of experts. Similarly, our investigation into the impact of MoAE layer count reveals that incorporating six MoAE layers yields the best performance, as shown in Fig. 3(b). Deviating from this number leads to a noticeable decline in performance. These findings suggest that replacing the final six ViT layers with MoAE layers establishes an optimal balance between task-specific adaptation and preservation of general knowledge.
Beyond structural parameters, we also explored the impact of external training data. As depicted in Fig. 3(c), increasing the proportion of the SFER dataset during the pre-training phase leads to consistent performance improvements on the DFEW dataset (up to 2.87%/3.04% UAR/WAR gain), whereas the impact on FERV39k is less pronounced (approximately 0.49%/1.07% UAR/WAR improvement). These differences may reflect variations in dataset characteristics, such as scale, label noise, or inherent task difficulty. Both datasets achieve peak performance when the entire SFER dataset is employed during pre-training, indicating that pre-training on SFER data is generally advantageous, especially for DFEW. Regarding the SFER dataset utilization during fine-tuning, our experiments reveal that increasing the proportion of SFER data from scratch enhances performance. Notably, using 50% of the SFER data yields optimal results, as illustrated in Fig. 3(d). However, further increasing the proportion beyond this point offers no additional benefits and could bias the optimization trajectory of the DFER task, potentially leading to a degradation in DFER performance.
Overall, these empirical findings highlight the importance of careful hyperparameter tuning in maximizing the effectiveness of the proposed unified learning framework. Specifically, using 8 MoAE experts, six MoAE layers, and the entire SFER dataset during pre-training, along with partial SFER data during fine-tuning, establishes a robust foundation for practical deployment while maintaining computational efficiency.
TABLE X: Ablation studies on the number of activated experts per MoAE layer.
Activated
Experts
FERV39K DFEW
UAR WAR UAR WAR
0 42.63 52.96 69.62 78.08
2 43.41 53.65 70.58 79.37
4 43.38 53.03 67.92 78.38
6 42.98 53.50 67.51 77.95
8 43.55 53.00 67.96 78.47
TABLE XI: Performance comparison of different curriculum strategies for incorporating SFER data during joint fine-tuning. Linear: gradually increasing the proportion of SFER data; Step: introducing SFER data only in the second half of training; Fixed: maintaining a constant 50% SFER proportion throughout.
Adjusting
Strategy
FERV39K DFEW
UAR WAR UAR WAR
Linear 41.78 52.66 66.52 76.76
Step 41.81 52.87 66.66 77.01
Fiexd (ours) 43.38 53.65 70.58 79.37
IV-D9 The Impact of Number Activated Experts
Ablation results in Table X demonstrate that activating 2 experts per MoAE layer yields the best overall performance, achieving the highest WAR on both FERV39K (53.65%) and DFEW (79.37%), as well as the highest UAR on DFEW (70.58%). Compared to the zero-expert setting, which essentially disables expert routing, the inclusion of even two experts significantly improves performance, confirming the functional value of our MoAE module. However, increasing the number of activated experts beyond 2 not only fails to improve performance but also introduces additional computational overhead, suggesting that excessive expert activation introduces redundancy and harms generalization. These findings indicate that a sparse expert activation strategy (i.e., top-2) provides the optimal balance between representational capacity and the risk of overfitting.
IV-D10 Comparison of SFER Data Adjusting Strategies
Ablation results in Table XI compare three curriculum strategies for incorporating SFER data into training. The proposed Fixed strategy, which maintains a constant 50% SFER proportion throughout training, outperforms both Linear (gradual increase) and Step (late introduction) baselines, achieving the best UAR/WAR on both FERV39K (43.38%/53.65%) and DFEW (70.58%/79.37%). Fixed integration allows the model to leverage semantic diversity from the outset, fostering better generalization and mitigating early overfitting to target-specific patterns, whereas delayed (Step) or gradual (Linear) incorporation limits the influence of auxiliary signals influence during the critical representation learning phase.
@@ IV-D9 The Impact of Number Activated Experts
TABLE XII: Cross-task evaluation comparisons on SFER (AffectNet-7) and DFER (FERV39K, DFEW).
@@
IV-E Correlation Analysis between SFER and DFER
IV-E1 Evaluation on Cross-Task
Since S4D can handle both SFER and DFER tasks using a shared backbone and respective classifiers, we treat it as a distinct task model during inference. To investigate the correlations and differences between the SFER and DFER tasks, we conducted cross-task evaluation comparisons on the AffectNet-7, FERV39K, and DFEW datasets. As shown in Table XII, the SFER model trained on AffectNet-7 achieves 32.34%/35.10% and 43.36%/47.13% UAR/WAR on the FERV39K and DFEW datasets, respectively. Although this performance is lower than that of DFER models on their respective tasks, it significantly exceeds random guessing (approximately 14.3%), indicating that the SFER model captures relevant spatial features from video data. In contrast, DFER models trained on the FERV39K and DFEW datasets achieve 51.46%/51.38% and 45.20%/45.04% UAR/WAR on AffectNet-7, respectively. These results suggest that DFER models can leverage static expression information from dynamic data, transferring well to the SFER task. We attribute this phenomenon to the inherent presence of static facial expressions within dynamic sequences, which capture peak expressions and serve as snapshots within the temporal progression of the video. However, direct comparisons of UAR/WAR metrics between SFER and DFER models are not ideal due to the inherent differences between the tasks. Therefore, these results should be viewed as reflecting feature transferability rather than a direct performance comparison.
@@ IV-E1 Evaluation on Cross-Task
Figure 5: Visualization of activation pathways. The figure shows the top-10 most frequently activated expert paths, where the top-2 are highlighted in color and the others in gray. Note that DFER and SFER tasks employ hard sharing through the FFN branch, enabling structural coupling while maintaining task-specific routing.
IV-E2 Semantic Relevance between SFER and DFER Tasks
@@
IV-E3 Expert Activation Patterns Analysis
To investigate the task-adaptive behavior of the proposed MoAE module, we visualize the most frequently activated expert paths across layers, as shown in Fig. 5. Although the DFER and SFER tasks share the FFN branch in the MoAE layer, the activated experts differ significantly across tasks. For instance, the expert selections in top-2 pathways for AffectNet-7 exhibit less overlap with those in FERV39K and DFEW, suggesting that the gating mechanism effectively preserves task-specific token routing and facilitates the learning of task-dependent representations. Notably, both DFER-related tasks (i.e., FERV39K and DFEW) and the SFER task (AffectNet-7) frequently activate expert E6 at layer L3, indicating the emergence of a shared mid-level expert. This observation implies that the model captures partially task-agnostic representations, thus enabling structural coupling beyond FFN sharing and promoting knowledge transfer across related tasks. Moreover, the activation paths in DFER exhibit greater consistency between the top-1 and top-2 selections, especially from layers L1 to L4, suggesting a more concentrated and deterministic routing pattern. In contrast, SFER demonstrates a more diverse expert selection strategy, relying on a larger number of expert activation pathways. These results indicate that MoAE simultaneously learns task-specific routing configurations while selectively reusing experts to encode cross-task shared features.
@@ IV-F2 Confusion Matrix Analysis
We performed a comprehensive statistical analysis of S4D’s performance on the DFEW [11], FERV39k [9], and MAFW [10] datasets, with the results presented through confusion matrices. As depicted in Fig. 7, majority categories such as happy, sad, and neutral expressions demonstrate high accuracy across all datasets. In contrast, minority categories, including disgust, fear, and helplessness exhibit lower accuracy, particularly on the DFEW dataset. This performance gap is likely attributable to the limited sample sizes and severe class imbalance in these categories. For example, disgust constitutes only 1.22% of the total samples in DFEW. Such challenges can be alleviated by implementing imbalanced learning strategies, such as oversampling techniques. Additionally, we observed that some expressions are frequently misclassified as neutral, particularly disgust and surprise, due to the subtle boundaries between neutral and these emotional expressions. For example, in both the DFEW and FERV39K datasets, disgust is frequently misclassified as neutral, likely because of the subtlety of disgust expressions, such as slight nose wrinkling or minor lip movements.